# Nearly-Linear Time Seeded Extractors with Short Seeds

### v1

@@ 5.1 A Non-Recursive Construction
In this section, we use the various previously introduced building blocks to construct a seeded extractor with order-optimal seed length $O(\log(n/\varepsilon))$ computable in time $\widetilde{O}(n)$. In a nutshell, our extractor proceeds as follows on input an $(n,k)$-source $X$:
@@
    Using a fresh seed, apply the lossless KT condenser from Theorem 3.10 to $X$. This yields an $(n^{\prime},k)$-source $X^{\prime}$ of length $n^{\prime}\approx k$ and constant entropy rate $\delta$ which can be arbitrarily close to $1$.
@@
    Using the fact that $X^{\prime}$ has high min-entropy rate, use the bounded-independence sampler from Lemma 3.9 to sample subsources from $X^{\prime}$ using a fresh seed. Specific properties of the bounded-independence sampler allow us to obtain a block source $Z=Z_{1}\circ Z_{2}\circ\cdots\circ Z_{t}$ with a seed of length only $O(\log(1/\varepsilon))$. The number of blocks is $t=O(\log n)$ and the blocks $Z_{i}$ have geometrically increasing lengths, up to an $n^{\alpha}$ length threshold.
@@
    Now, to prepare for the hash-based iterative extraction, we need to make our blocks decreasing. Again, using a short seed, of length $O(\log(n/\varepsilon))$, we transform $Z$ into $S=S_{1}\circ\cdots S_{t}$, where the blocks are now geometrically decreasing. The blocks lengths will vary from $n^{\beta_{1}}$ to some $n^{\beta_{2}}$, for some constants $\beta_{1}>\beta_{2}$.
@@
    Using a fresh seed, apply the fast hash-based extractor from Lemma 2.14 to perform block source extraction from $S$. Noting that the first block has length $n^{\Omega(1)}$, the block source extraction only outputs $n^{\Omega(1)}$ bits. We are able to use only $O(\log(n/\varepsilon))$ random bits here, since we do not output $n^{\Omega(1)}$ bits already at the beginning of the iterative extraction process, but instead output logarithmically many bits, and gradually increasing the output length.
These steps will culminate in the following theorem.
@@
There exists a constant $c\in(0,1)$ such that for every positive integers $n$ and $k\leq n$, any $\varepsilon\geq 2^{-k^{c}}$, and any constant $\eta\in(0,1)$, there exists a strong $(k,\varepsilon)$ extractor
@@
5.1.1 Item 2: Generating the block source
Because of the initial condensing step, we will assume from here onwards that our input source $X$ is an $(n,k=\delta n)$-source with constant $\delta$. In order to generate the desired block source, we first use a fresh seed $Y$ as input to an appropriate instantiation of the bounded-independence sampler $\Gamma$ from Lemma 3.9. This yields a tuple of coordinates $\Gamma(Y)=j_{1},\dots,j_{m_{t}}$ from $[n]$, such that $\Gamma(Y)|_{[1,m_{i}]}$ is an appropriate averaging sampler for every $i$. Then, we use these coordinates to sample subsources from $X\sim\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}$, and get a block source with increasing blocks.
@@ 5.1.1 Item 2: Generating the block source
  • A closeness parameter $\varepsilon\in(0,1)$ that satisfies $\varepsilon\geq 2^{-c_{\varepsilon}n}$ where $c_{\varepsilon}=c_{\varepsilon}(\zeta,\delta)$ is constant,
@@
takes an independent and uniform random seed $Y\sim\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d_{\mathsf{samp}}}$ and outputs a random variable $Z$ that is $\varepsilon$-close to a
@@
block-source, where each $\Delta_{i-1}=\alpha\cdot\Delta_{i}$ for $\alpha=\frac{\zeta\delta}{4}$. Moreover, the seed length $d=O\mathopen{}\mathclose{{}\left(\frac{\log n}{\log\Delta_{1}}\cdot\log\frac{t}{\varepsilon}}\right)$, and the procedure runs in time $\widetilde{O}(n+\log^{2}(t/\varepsilon))$.
Note that for any constants $10<\theta_{1}<\theta_{2}<$ 1, and any $\varepsilon=\Omega(2^{-\sqrt{n}})$, we can have $\Delta_{t}=n^{\theta_{2}}$ and $\Delta_{1}=n^{\theta_{1}}$ for some $t=O(\log n)$, with seed length $O(\log(1/\varepsilon))$ and runtime $\widetilde{O}(n)$.
@@
by choosing the constant $c_{\mathsf{t}}$ appropriately. Let $\Gamma\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d_{\mathsf{samp}}}\rightarrow[n]^{m_{t}}$ be the $(\gamma,\varepsilon_{\Gamma})$ sampler of Lemma 3.9, set with $\varepsilon_{\Gamma}=\frac{1}{\log(\frac{6}{\zeta\delta})}\cdot\frac{\zeta\delta}{6}=O(1)$ and $\gamma=\frac{\varepsilon}{2t}$. Note that then,
@@
and indeed $m_{t}\leq\frac{\varepsilon_{\Gamma}}{16}\cdot n$ can be met by, again, setting the constant $c_{\mathsf{t}}$ appropriately. Moreover, we have that for any $i\in[t]$,
@@
is a $(\gamma,\varepsilon_{\Gamma})$ sampler, where $w\sim W_{i}$ has distinct symbols. Set $\beta=\frac{\zeta}{2}$.
@@
Now, Lemma 2.16, instantiated with $\tau=\frac{\beta\delta}{3}$ (notice that indeed $\varepsilon_{\Gamma}\leq\frac{\tau}{\log(1/\tau)}$), tells us that for every $i\in[t]$, denoting $A_{i}=X_{W_{i}}$, there exists a set $\mathbf{B}_{i}\subseteq\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d_{\mathsf{samp}}}$ of bad $y$-s of density at most $\gamma+2^{-\Omega(\tau n)}$, such that for any $y\notin\mathbf{B}_{i}$,
$A_{i}|\mathopen{}\mathclose{{}\left\{{Y=y}}\right\}\sim\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m_{i}}$
has entropy rate $\delta-3\tau\geq(1-\beta)\delta$ for every $y$. Letting $Z=A_{t}$, union-bounding over the bad $y$-s tells us that $Z$ is $t\cdot(\gamma+2^{-\Omega(n)})\leq\varepsilon$ close to some $Z^{\prime}\sim\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m_{t}}$ such that for any $i\in[t]$, $Z^{\prime}_{i}=Z^{\prime}_{[1,m_{i}]}\sim\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m_{i}}$ has entropy rate $(1-\beta)\delta$.
Next, we apply the chain rule for min-entropy to argue that $Z^{\prime}$ (and hence $Z$) is close to a block source. To do that, we apply the chain rule for min-entropy $t-1$ times. For simplicity, abbreviate $Z^{(i)}=Z^{\prime}_{[m_{i-1}+1,m_{i}]}$ (so note that $Z^{\prime}_{i}$ is the longer block, $Z^{\prime}_{[1,m_{i}]}$, whereas $Z^{(i)}$ is its length-$\Delta_{i}$ suffix), so
$Z^{\prime}=\mathopen{}\mathclose{{}\left(Z^{(1)},Z^{(2)},\ldots,Z^{(t)}}\right).$
We will argue that $Z^{\prime}$ is a block source. Applying Lemma 2.5, we know that for any $i\in[t]$,
@@
$\widetilde{\mathbf{H}}_{\infty}\mathopen{}\mathclose{{}\left(Z^{(i)}\mid Z^{(1)},\ldots,Z^{(i-1)}}\right)\geq\mathbf{H}_{\infty}(Z^{(i)})-\sum_{j=1}^{i-1}\Delta_{j}=\mathbf{H}_{\infty}(Z^{(i)})-m_{i-1}\geq\mathbf{H}_{\infty}(Z^{(i)})-\frac{\alpha}{1-\alpha}\Delta_{i}.$
Now, $Z^{\prime}_{i}=(Z^{\prime}_{i-1},Z^{(i)})$, so $\mathbf{H}_{\infty}(Z^{(i)})\geq\mathbf{H}_{\infty}(Z^{\prime}_{i})-m_{i-1}$, and notice that
$(1-\beta)\delta\cdot\Delta_{i}-m_{i-1}\geq(1-\beta)\delta\cdot\Delta_{i}-\frac{\alpha}{1-\alpha}\Delta_{i}\geq(1-\zeta)\delta\cdot\Delta_{i},$
where we used the fact that $\alpha\leq\frac{(\zeta-\beta)\delta}{1-(\zeta-\beta)\delta}$.
@@
The bound on the runtime follows easily, recalling that $\Gamma$ runs in time $\widetilde{O}\mathopen{}\mathclose{{}\left(n+\log^{2}(1/\gamma)}\right)$. ∎
5.1.2 Item 3: Subsampling from the block source
@@
To apply iterative extraction, we will our block source to have decreasing blocks. Here, we will use a sampler to sample from each block, using the same seed accross the blocks.
@@
block-source $Z=(Z_{1},\ldots,Z_{t})$, for every $\Delta_{1}\leq\ldots\leq\Delta_{t}$ and a constant $\delta$, and:
@@
takes an independent and uniform random seed $Y\sim\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d_{\mathsf{samp}}}$ and outputs a random variable $S$ that is $\varepsilon$-close to a $\mathopen{}\mathclose{{}\left((\ell_{1},\ldots,\ell_{t}),(1-\zeta)\delta}\right)$ block-source, where each $\ell_{i+1}=\alpha\cdot\ell_{i}$, and assuming that $\ell_{t}\geq c_{1}\log(t/\varepsilon)$ where $c_{1}=c_{1}(\zeta,\delta)$ is a constant. Moreover, the seed length $d=\log\frac{\Delta_{t}}{\ell_{1}}+O\mathopen{}\mathclose{{}\left(t+\log\frac{1}{\varepsilon}}\right)$, and the procedure runs in time $\operatorname{polylog}(\Delta_{t})\cdot\ell_{1}$.
Note that when $\Delta_{1}=n^{\theta_{1}}$ and $\ell_{t}=n^{\beta}$ for some constants $20<\beta<\theta_{1}<$ 2, $d_{\mathsf{samp}}=O(\log(n/\varepsilon))$, the procedure runs in time $O(n)$, and we can take any $\varepsilon\geq 2^{-c\cdot\ell_{t}}$ for some constant $c$ that depends on $\zeta$ and $\delta$.
@@
For $i\in[t]$, let $m_{i}=\sum_{j=1}^{i}\ell_{i}$, recalling that $\ell_{i}=\alpha^{i-1}\ell_{1}$. For each $i\in[t]$, let $\Gamma_{i}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d_{i}}\rightarrow[\Delta_{i}]^{\ell_{i}}$ be the $(\gamma,\varepsilon_{\Gamma})$ be the distinct-samplers sampler of Lemma 2.21, where $\gamma=\frac{\varepsilon}{2t}$ and $\varepsilon_{\Gamma}=\frac{1}{\log(\frac{6}{\zeta\delta})}\cdot\frac{\zeta\delta}{6}=O(1)$. We need to make sure that each $\ell_{i}\geq c\cdot\frac{\log(1/\gamma)}{\varepsilon^{2}_{\Gamma}}$ for some universal constant $c$, and indeed that is the case, by our constraint on $\ell_{t}$. Also, $d_{i}=\log(\Delta_{i}/\ell_{i})+O(\log\frac{1}{\gamma}\cdot\operatorname{poly}(1/\varepsilon_{\Gamma}))$ and we set $d_{\mathsf{samp}}$ to be the maximal over the $d_{i}$-s, so

### v2

@@ 4.1 A Non-Recursive Construction
In this section, we use the sampler based on bounded independence from section˜3.2 and the nearly-linear time KT condenser from section˜3.3 to construct a seeded extractor with order-optimal seed length $O(\log(n/\varepsilon))$ computable in time $\widetilde{O}(n)$. We remark that the goal of this section is to give a low-error, relatively simple (and in particular, non-recursive) construction. Among the non-recursive “sample-then-extract” extractor constructions, two notable ones are [NZ96] and [ZUC96]: The [NZ96] construction uses poly-logarithmic seed (see footnote˜26); The [ZUC96] construction bears some resemblance to our construction and also utilizes sub-sampling, but only works in the high-error regime, namely has seed length $\Theta(\varepsilon^{-2}+\log n)$. Constructions in that framework that support low error and short seed include [SZ99] and followup works such as [GUV09, ZUC97]. We adapt those recursive constructions in section˜4.2 to get our nearly linear time construction of ˜1. Finally, we note that other block source conversion techniques have been used for constructing pseudorandomness primitives, such as in [TA-02, DKS+13], but they are less relevant in our context.
In a nutshell, our extractor proceeds as follows on input an $(n,k)$-source $X$:
@@
    Using a fresh seed, apply the lossless KT condenser from theorem˜3.14 to $X$. This yields an $(n^{\prime},k)$-source $X^{\prime}$ of length $n^{\prime}\approx k$ and constant entropy rate $\delta$ which can be arbitrarily close to $1$. Note that in the parameter regime considered in this section the KT condenser does not require the one-time preprocessing step.
@@
    Using the fact that $X^{\prime}$ has high min-entropy rate, use the bounded-independence sampler from lemma˜3.9 to sample subsources from $X^{\prime}$ using a fresh seed. Specific properties of the bounded-independence sampler allow us to obtain a block source $Z=(Z_{1},Z_{2},\dots,Z_{t})$ with a seed of length only $O(\log(1/\varepsilon))$. The number of blocks is $t=O(\log n)$ and the blocks $Z_{i}$ have geometrically increasing lengths, up to an $n^{\alpha}$ length threshold.
@@
    Now, to prepare for the hash-based iterative extraction, we need to make our blocks decreasing. Again, using a short seed, of length $O(\log(n/\varepsilon))$, we transform $Z$ into $S=(S_{1},\dots,S_{t})$, where the blocks are now geometrically decreasing. The blocks lengths will vary from $n^{\beta_{1}}$ to some $n^{\beta_{2}}$, for some constants $\beta_{1}>\beta_{2}$. (Here, we do not use the “prefix samplers” property of lemma˜3.9, so transforming $Z$ into $S$ can be done using an existing sampler, and specifically the one from lemma˜2.20.)
@@
    Using a fresh seed, apply the fast hash-based extractor from lemma˜2.13 to perform block source extraction from $S$. Noting that the first block has length $n^{\Omega(1)}$, the block source extraction only outputs $n^{\Omega(1)}$ bits.
    While the seed length of lemma˜2.13 requires at least $m$ random bits, we are still able to use only $O(\log(n/\varepsilon))$ bits, since we do not output $n^{\Omega(1)}$ bits already at the beginning of the iterative extraction process, but instead first output logarithmically many bits, and then gradually increase the output length.
Finally, once we have extracted $n^{\Omega(1)}$ random bits, outputting almost all the entropy can be done using standard techniques (see section˜4.1.4). These steps will culminate in the following theorem.
@@
There exists a constant $c\in(0,1)$ such that for every positive integers $n$ and $k\leq n$, any $2^{-k^{c}}\leq\varepsilon\leq\frac{1}{n}$, and any constant $\eta\in(0,1)$, there exists a strong $(k,\varepsilon)$ extractor
@@
4.1.1 item˜2: Generating the block source
Because of the initial condensing step, we will assume from here onwards that our input source $X$ is an $(n,k=\delta n)$-source with constant $\delta$. In order to generate the desired block source, we first use a fresh seed $Y$ as input to an appropriate instantiation of the bounded-independence sampler $\Gamma$ from lemma˜3.9. This yields a tuple of coordinates $\Gamma(Y)=j_{1},\dots,j_{m_{t}}$ from $[n]$, such that $\Gamma(Y)|_{[1,m_{i}]}$ is an appropriate averaging sampler for every $i$. Then, we use these coordinates to sample subsources from $X\sim\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}$, and get a block source with increasing blocks. We recall that getting increasing blocks is only an intermediate step towards our final goal. Indeed, in this step, we sample from the source $X$, which will guarantee that a typical prefix leaves enough entropy in the next, larger, blocks. In section˜4.1.2 we will need to subsample from those blocks and argue that each block still has entropy even after a typical fixing of all the blocks that precedes it. The latter property is the one needed for block-source extraction.
@@
  • A closeness parameter $\varepsilon\in(0,1)$,
@@
takes an independent and uniform random seed $Y\sim\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d_{\mathsf{samp}}}$ and outputs a random variable $Z$ such that, for every $y$, $(Z|Y=y)$ is $\varepsilon$-close to an exact
@@
block-source, where each $\Delta_{i-1}=\alpha\cdot\Delta_{i}$ for $\alpha=\frac{\zeta\delta}{16}$ and assuming that $\varepsilon\geq 2^{-(c_{\varepsilon}\Delta_{1}-t)}$ for some constant $c_{\varepsilon}=c_{\varepsilon}(\zeta,\delta)$. Moreover, the seed length $d=O\mathopen{}\mathclose{{\left(\frac{\log n}{\log\Delta_{1}}\cdot\log\frac{t}{\varepsilon}}}\right)$, and the procedure runs in time $\widetilde{O}(n)+\operatorname{polylog}(t/\varepsilon)$.
Note that for any constants $0<\theta_{1}<\theta_{2}<1$, and any $\varepsilon\geq 2^{-n^{c}}$ where $c>0$ is a small enough constant, we can have $\Delta_{t}=n^{\theta_{2}}$ and $\Delta_{1}=n^{\theta_{1}}$ for some $t=O(\log n)$, with seed length $O(\log(t/\varepsilon))$ and runtime $\widetilde{O}(n)$.
@@
by choosing the constant $c_{\mathsf{t}}$ appropriately. Let $\Gamma\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d_{\mathsf{samp}}}\rightarrow[n]^{m_{t}}$ be the $(\gamma,\varepsilon_{\Gamma})$-averaging sampler of lemma˜3.9, set with $\varepsilon_{\Gamma}=\frac{1}{\log(\frac{6}{\zeta\delta})}\cdot\frac{\zeta\delta}{6}=O(1)$ and $\gamma=\frac{\varepsilon}{8t}$. Note that then,
@@
and indeed $m_{t}\leq\frac{\varepsilon_{\Gamma}}{16}\cdot n$ can be met by, again, setting the constant $c_{\mathsf{t}}\in(0,1)$ appropriately. Moreover, we have that for any $i\in[t]$,
@@
is a $(\gamma,\varepsilon_{\Gamma})$ sampling distribution, where $w\sim W_{i}$ has distinct symbols. Set $\beta=\frac{\zeta}{2}$.
For each $i\in[t]$, let $A_{i}=X_{W_{i}}$. We may write $A_{t}=(Z_{1},\dots,Z_{t})$ with $Z_{j}=(A_{t})_{[m_{j-1}+1,m_{j}]}$. Note that under this perspective we have $A_{i}=(Z_{1},\dots,Z_{i})$ for each $i\in[t]$. We claim that $(Z_{1},\dots,Z_{t})$ is close to an appropriate (exact) block source $Z^{\prime}=(Z^{\prime}_{1},\dots,Z^{\prime}_{t})$, even conditioned on the seed $Y$. This follows by an induction argument similar to that of [NZ96, proof of Lemma 17] which we detail below.
@@
First, lemma˜2.15 instantiated with $\tau=\frac{\beta\delta}{3}$ (notice that indeed $\varepsilon_{\Gamma}\leq\frac{\tau}{\log(1/\tau)}$) tells us that
$(Y,Z_{1},\dots,Z_{i})=(Y,A_{i})\approx_{\gamma+2^{-\Omega(\tau n)}}(Y,A^{\prime}_{i}),$ (1)
where $(A^{\prime}_{i}|Y=y)$ has entropy rate at least $\delta-3\tau\geq(1-\beta)\delta$ for every $y$. eq.˜1 with $i=1$ and $Z^{\prime}_{1}=A^{\prime}_{1}$ is our base case. Now, fix an arbitrary $i\geq 2$ and suppose that we already know that
$(Y,Z_{1},\dots,Z_{i-1})\approx_{2(i-1)(\gamma+2^{-\Omega(\tau n)}+\xi)}(Y,Z^{\prime}_{1},\dots,Z^{\prime}_{i-1}),$ (2)
where $(Z^{\prime}_{j}|Y=y,Z^{\prime}_{1}=z_{1},\dots,Z^{\prime}_{j-1}=z_{j-1})$ has entropy rate at least $(1-\zeta)\delta$ for every $1\leq j\leq i-1$ and every $y,z_{1},\dots,z_{j-1}$, and $\xi=\frac{\varepsilon}{4t}$.
Write $A^{\prime}_{i}=(Z^{\prime\prime}_{1},\dots,Z^{\prime\prime}_{i})$ with $|Z^{\prime\prime}_{j}|=|Z^{\prime}_{j}|$. Applying lemma˜2.6 to $(A^{\prime}_{i}|Y=y)$ with $\delta=\xi$ yields
$\mathbf{H}_{\infty}(Z^{\prime\prime}_{i}|Y=y,Z^{\prime\prime}_{1}=z_{1},\dots,Z^{\prime\prime}_{i-1}=z_{i-1})$ $=\mathbf{H}_{\infty}(A^{\prime}_{i}|Y=y,Z^{\prime\prime}_{1}=z_{1},\dots,Z^{\prime\prime}_{i-1}=z_{i-1})$
$\geq\mathbf{H}_{\infty}(A^{\prime}_{i})-\sum_{j=1}^{i-1}\Delta_{j}-\log(1/\xi)$
$=\mathbf{H}_{\infty}(A^{\prime}_{i})-m_{i-1}-\log(1/\xi)$
$\geq(1-\beta)\delta m_{i}-\frac{\alpha}{1-\alpha}\Delta_{i}-\log(1/\xi)$ (3)
except with probability at most $\xi$ over the choice of $(z_{1},\dots,z_{i-1})$. Using the fact that $m_{i}\geq\Delta_{i}$, to get entropy rate at least $(1-\zeta)\delta$ it is left to verify that
@@
$\mathopen{}\mathclose{{\left((\zeta-\beta)\delta-\frac{\alpha}{1-\alpha}}}\right)\Delta_{i}=\mathopen{}\mathclose{{\left(\frac{\beta\delta}{2}-\frac{\alpha}{1-\alpha}}}\right)\Delta_{i}\geq\log(1/\xi).$
Using our bound on $\alpha$, we get that $\frac{\alpha}{1-\alpha}\leq\frac{\beta\delta}{4}$. Thus, $\frac{\beta\delta}{2}\Delta_{i}\geq\log(1/\xi)$ holds whenever $\log(1/\varepsilon)\leq c_{\varepsilon}\Delta_{1}-t$, where both $c_{\varepsilon}$ depend only on $\delta$ and $\zeta$.
Call a vector $\vec{v}=(y,z_{1},\dots,z_{i-1})$ good if eq.˜3 holds. Suppose that we already have blocks $B_{1},\dots,B_{i-1}$, arbitrarily distributed. We generate one more block $B_{i}$ as follows. First, sample $\vec{v}\sim(Y,B_{1},\dots,B_{i-1})$. If $\vec{v}$ is good as defined above, we set $B_{i,\vec{v}}$ (the random variable $B_{i}$ conditioned on $(Y,B_{1},\dots,B_{i-1})=\vec{v}$) to be $B_{i,\vec{v}}=(Z^{\prime\prime}_{i}|(Y,Z^{\prime\prime}_{1},\dots,Z^{\prime\prime}_{i-1})=\vec{v})$. Otherwise, including when $\vec{v}$ is not in the support of $(Y,Z^{\prime\prime}_{0},\dots,Z^{\prime\prime}_{i-1})$, we set $B_{i,\vec{v}}$ to be uniformly distributed over $\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{\Delta_{i}}$ and independent of everything else. Note that by construction we have $\mathbf{H}_{\infty}(B_{i,\vec{v}})\geq(1-\zeta)\delta\Delta_{i}$ for all $\vec{v}$.
Then, it follows from eq.˜1 that
$\mathopen{}\mathclose{{\left(Y,Z_{1},\dots,Z_{i-1},B^{(1)}_{i}}}\right)\approx_{\gamma+2^{-\Omega(\tau n)}}\mathopen{}\mathclose{{\left(Y,Z^{\prime\prime}_{1},\dots,Z^{\prime\prime}_{i-1},B^{(2)}_{i}}}\right)$, (4)
where on the left-hand side we take $B^{(1)}_{i}$ to be sampled based on $Y$ and $B_{j}=Z_{j}$ for $1\leq j\leq i-1$ as described for $B_{i}$ in the previous paragraph, and on the right-hand side we take $B^{(2)}_{i}$ to be sampled based on $Y$ and $B_{j}=Z^{\prime\prime}_{j}$ for $0\leq j\leq i-1$. Since $\vec{v}\sim(Y,Z^{\prime\prime}_{1},\dots,Z^{\prime\prime}_{i-1})$ is good with probability at least $1-\xi$, in which case $B_{i,\vec{w}}$ is sampled identically to $Z^{\prime\prime}_{i,\vec{w}}$, we get from eq.˜4 and a triangle inequality that
$\mathopen{}\mathclose{{\left(Y,Z_{1},\dots,Z_{i-1},B^{(1)}_{i}}}\right)\approx_{\gamma+2^{-\Omega(\tau n)}+\xi}\mathopen{}\mathclose{{\left(Y,Z^{\prime\prime}_{1},\dots,Z^{\prime\prime}_{i-1},Z^{\prime\prime}_{i}}}\right).$ (5)
By eq.˜2, we also have that
$\mathopen{}\mathclose{{\left(Y,Z_{1},\dots,Z_{i-1},B^{(1)}_{i}}}\right)\approx_{2(i-1)(\gamma+2^{-\Omega(\tau n)}+\xi)}\mathopen{}\mathclose{{\left(Y,Z^{\prime}_{1},\dots,Z^{\prime}_{i-1},B_{i}^{(3)}}}\right),$ (6)
where, again, $B_{i}^{(3)}$ is sampled based on $Y$ and $B_{j}=Z^{\prime}_{j}$ for $1\leq j\leq i-1$ as described for $B_{i}$ above. Combining eq.˜1 (recall that $A^{\prime}_{i}=(Z^{\prime\prime}_{1},\dots,Z^{\prime\prime}_{i})$) with eqs.˜5 and 6 via a triangle inequality, we get that
$\mathopen{}\mathclose{{\left(Y,Z_{1},\dots,Z_{i}}}\right)\approx_{2i(\gamma+2^{-\Omega(\tau n)}+\xi)}\mathopen{}\mathclose{{\left(Y,Z^{\prime}_{1},\dots,Z^{\prime}_{i-1},B^{(3)}_{i}}}\right)$.
Note that the sampling of $B^{(3)}_{i}$ on the right-hand side of this equation guarantees that $\mathbf{H}_{\infty}(B^{(3)}_{i}|Y=y,Z^{\prime}_{1}=z_{1},\dots,Z^{\prime}_{i-1}=z_{i-1})\geq(1-\zeta)\delta\Delta_{i}$ for all $(y,z_{1},\dots,z_{i-1})$. Therefore, $(Z^{\prime}_{1},\dots,Z^{\prime}_{i-1},Z^{\prime}_{i}=B^{(3)}_{i})$ is indeed the target block source with $i$ blocks. Setting $i=t$, and by inspection of the sampling process for the $Z^{\prime}_{j}$-s, gives that $(Z|Y=y)$ is
$2t\cdot\mathopen{}\mathclose{{\left(\gamma+2^{-\Omega(\tau n)}+\xi}}\right)\leq\varepsilon$
close to an exact $((\Delta_{1},\ldots,\Delta_{t}),(1-\zeta)\delta)$ block source for every $y$.
@@
The bound on the runtime follows easily, recalling that $\Gamma$ runs in time $\widetilde{O}(n)+\operatorname{polylog}(1/\gamma)$. ∎
4.1.2 item˜3: Subsampling from the block source
@@
To apply iterative extraction, we will need our block source to have decreasing blocks. Here, we will use a sampler to sample from each block, using the same seed across the blocks.
@@
block-source $Z=(Z_{1},\ldots,Z_{t})$, where $\Delta_{1}\leq\ldots\leq\Delta_{t}$ and $\delta$ is a constant,
@@
  • An independent and uniform random seed $Y\sim\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d_{\mathsf{samp}}}$,
satisfies the following. Assuming that $\ell_{t}\geq c_{1}\log(t/\varepsilon)$ for some constant $c_{1}=c_{1}(\zeta,\delta)$, it outputs a random variable $S$ such that, for every $y$, $(S|Y=y)$ is $\varepsilon$-close to an exact
$\mathopen{}\mathclose{{\left((\ell_{1},\ldots,\ell_{t}),(1-\zeta)\delta}}\right)$
block-source, where each $\ell_{i+1}=\alpha\cdot\ell_{i}$ Moreover, the seed length $d_{\mathsf{samp}}=\log\frac{\Delta_{t}}{\ell_{1}}+O\mathopen{}\mathclose{{\left(t+\log\frac{1}{\varepsilon}}}\right)$, and the procedure runs in time $t\cdot\operatorname{polylog}(\Delta_{t})\mathopen{}\mathclose{{\left(\ell_{1}+\log^{2}(t/\varepsilon)}}\right)$.
Note that when $\Delta_{t}=n^{\theta_{1}}$ and $\ell_{1}=n^{\beta}$ for some constants $\theta_{1},\beta\in(0,1)$, $d_{\mathsf{samp}}=O(\log(n/\varepsilon))$, the procedure runs in time $O(n)$, and we can take any $\varepsilon\geq 2^{-c\cdot\ell_{t}}$ for some constant $c$ that depends on $\zeta$ and $\delta$.
@@
For $i\in[t]$, let $m_{i}=\sum_{j=1}^{i}\ell_{i}$, recalling that $\ell_{i}=\alpha^{i-1}\ell_{1}$. For each $i\in[t]$, let $\Gamma_{i}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d_{i}}\rightarrow[\Delta_{i}]^{\ell_{i}}$ be the $(\gamma,\varepsilon_{\Gamma})$ distinct-samples sampler of lemma˜2.20, where $\gamma=\frac{\varepsilon}{2t}$ and $\varepsilon_{\Gamma}=\frac{1}{\log(\frac{6}{\zeta\delta})}\cdot\frac{\zeta\delta}{6}=O(1)$. We need to make sure that each $\ell_{i}\geq c\cdot\frac{\log(1/\gamma)}{\varepsilon^{2}_{\Gamma}}$ for some universal constant $c$, and indeed that is the case, by our constraint on $\ell_{t}$. Also, $d_{i}=\log(\Delta_{i}/\ell_{i})+O(\log\frac{1}{\gamma}\cdot\operatorname{poly}(1/\varepsilon_{\Gamma}))$ and we set $d_{\mathsf{samp}}$ to be the maximum over the $d_{i}$-s, so